The density of a gas is 2.08 g/Lg/L at a temperature of 21 ∘C∘C and a pressure of 0.784 atmatm . Calculate the molar mass of the gas.
Answer #1

density of gas is 2.08 g/L means mass = 2.08 g and volume= 1 L

now using

PV= nRT

P= pressure = 0.784atm     

V= volume = 1L

n= number of moles =?

R= gas constant = 0.0821 LatmmolK

T = temperature = 21+273.15= 294.15 K

n= PV/RT= (0.784*1)/(0.0821*294.15)= 0.03246 moles

molar mass of gas= mass/moles = 2.08 g/ 0.03246 mol = 64 g/mol